OpenStudy (anonymous):

find y' if 3xy^2 + y^3=e^(x+y)

OpenStudy (anonymous):

\[3y^2+6xyy'+3y^2y'=e^{x+y}+y'e^{(x+y)}

OpenStudy (anonymous):

\[3y^2+6xyy'+3y^2y'=e^{x+y}+y'e^{(x+y)}\]

OpenStudy (anonymous):

first term required the produce rule rhs required chain rule

OpenStudy (anonymous):

so far so good or no?

OpenStudy (anonymous):

because it is algebra from here on in. solve for y'

OpenStudy (anonymous):

algebra no good under sleep deprivation, lol :] but yeah, so far makes sense

OpenStudy (anonymous):

\[6xyy'+3y^2y'-y'e^{(x+y)}=e^{(x+y)}-3y^2\] \[y'(6xy+3y^2-e^{(x+y)})=e^{(x+y)}-3y^2\] \[y'=\frac{e^{(x+y)}-3y^2}{6xy+3y^2-e^{(x+y)}}\]

OpenStudy (anonymous):

no guarantees on the algebra

OpenStudy (anonymous):

It actually looks good, and I'm sure that's a lot more accurate than I would've produced now. Thanks again, man. Much appreciated!